OpenStudy (anonymous):

What is the slope of the line between (–4, 4) and (–1, –2)? 1 2 –2 –1

OpenStudy (anonymous):

\[slope = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ (-2) - 4 }{ (-1) - (-4) }\]

OpenStudy (anonymous):

-6/3

OpenStudy (anonymous):

so -2?

OpenStudy (anonymous):

am i right ?

OpenStudy (anonymous):

Yep (:

Parth (parthkohli):

Heh, just messing, you're right. ;)